OpenStudy (archtangent):

my integrating skills are getting rusty :( what is the integral of y/x?? or is that even posible? ive checked an online integrating calculator and... ∫f(x,y)dx= ln(x)⋅y this is the result? can anyone confirm??

OpenStudy (anonymous):

Isn't that y will be treated as constant, when your integrand is dx?

OpenStudy (jhannybean):

Yeah.

OpenStudy (jhannybean):

\[\int \frac{y}{x}dx\] Take out y as a constant, because you are integrating with respect to x. \[y\int \frac{1}{x}dx\]

OpenStudy (anonymous):

\[\int\limits \frac{y}{x} \cdot dx \implies y \cdot \int\limits \frac{1}{x} \cdot dx \implies y \cdot \ln(x)\]

OpenStudy (anonymous):

+ C as well..

OpenStudy (jhannybean):

\[\dfrac{d}{dx} ln(x) = \frac{1}{x} \]\[\int \frac{1}{x}dx = \ln(x) +c\]

OpenStudy (archtangent):

ahhh yes, shame i overlooked that, yes thanks for pointing it out good job, thumbs up thanks maam's :))

OpenStudy (anonymous):

There is only one Ma'am here. :P

OpenStudy (anonymous):

Anyways, you are welcome.. :)

OpenStudy (archtangent):

thank you peeps then ehehe :))

OpenStudy (perl):

you need to add a function h(y) to your integral

OpenStudy (archtangent):

yup, noted, thanks sir thumbs up

OpenStudy (perl):

since with respect to x, the derivative of h(y) is zero

OpenStudy (zzr0ck3r):

Is there something like implicit integration @perl